San Diego Sheet Metal, Inc. incurs a variable cost of $40 per pound for raw material to produce a special alloy used in manufacturing aircraft. Required: In the table below, show the unit cost and total cost of raw material at the following production levels: 1 pound, 10 pounds, and 1,009 pounds. Production Level in pounds Unit Cost Total Cost per pound per pound per pound 10 1,000 Prev 17 of 19 l Next 28 29 30
Answer #1

When production level is 1 pound:

Unit cost= $40 per pound

Total cost= 40*1= $40

When production level is 10 pounds:

Unit cost= $40 per pound

Total cost= 40*10= $400

When production level is 1000 pounds:

Unit cost= $40 per pound

Total cost= 40*1000= $40000